A billboard is 27 feet high and casts a 36 foot shadow at noon. At the same time, a fir tree next to the billboard casts a 48 foot shadow. What proportion can be used to find the height, H, of the fir tree?
Found 2 solutions by edjones, mananth:
Answer by edjones(8007)
Answer by mananth(16949)

shadow of billboard = 36 feet
height = 27 feet
..
shadow of fir tree = 48 feet
height = x
..
Direct proportion
36/27 = 48/x
x= 48*27/36
x=36 feet the height of fir tree